OpenStudy (anonymous):

Point -Slope Form y-b=m(x-a) (x,y)(a,b) Question (5,5) (3,2)

OpenStudy (shadowlegendx):

So what is the question?

OpenStudy (anonymous):

how do you find the slope and write the equation in point-slope form.

OpenStudy (shadowlegendx):

Texas can probably explain this process in great detail :P

OpenStudy (anonymous):

ok thanks

OpenStudy (texaschic101):

find the slope using the slope form : slope(m) = (y2 - y1) / (x2 - x1) (5,5) x1 = 5 and y1 = 5 (3,2) x2 = 3 and y2 = 2 now we sub slope(m) = (2 - 5) / (3 - 5) slope (m) = -3/-2 slope(m) = 3/2 you with me so far ?

OpenStudy (texaschic101):

any questions so far ?

OpenStudy (anonymous):

thanks you so much :)

OpenStudy (texaschic101):

now we need to put it in point slope form....correct ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

point slope form = y - y1 = m(x - x1) slope(m) = 3/2 you can choose either of your points....(5,5) x1 = 5 and y1 = 5 now we sub y - 5 = 3/2(x - 5) ==> point slope form

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

i still don't understand.

OpenStudy (texaschic101):

the slope or the point slope form ?

OpenStudy (anonymous):

both

OpenStudy (texaschic101):

these are just formulas you will be using... slope(m) : (y2 - y1) / (x2 - x1) slope intercept form : y = mx + b point slope form : y - y1 = m(x - x1) you will use your 2 set of points ...(x1,y1) and (x2,y2). Once you have your slope(m), just sub that , along with 1 set of points, into the point slope form. It is basically all just a matter of subbing stuff in.

OpenStudy (anonymous):

what do we put in (y2 - y1) / (x2 - x1)

OpenStudy (texaschic101):

that is the formula to find the slope. You do that first.

OpenStudy (texaschic101):

because you need the slope to be able to sub it into the point slope formula

OpenStudy (texaschic101):

oh....I see what you are saying..... we use the set of points and put them into the slope formula... first set (x1,y1) and second set (x2,y2)

OpenStudy (anonymous):

what number go in those number

OpenStudy (anonymous):